# EDEN: Exploring Disks Embedded in N-body simulations of Milky-Way-mass halos from Symphony

### v1

@@ 3.1 Background
The language that the field uses to describe mass loss in subhalos paints a picture in which subhalos are completely destroyed in a cataclysmic, discrete event: subhalos are said to “disrupt.” In previous embedded disk simulations (Garrison-Kimmel et al., 2017; Kelley et al., 2019), the abundance suppression in the subhalo peak mass function (SPMF) is used as a diagnostic for ‘disruption’ of subhalos instead of enhanced mass loss. The peak mass of subhalos do not change after infall, so if the subhalos only experience enhanced stripping and could still be tracked by the halo finder, there would be no suppression in the SPMF in the ideal case; whereas ‘disruption’ of subhalos would lead to objects no longer detectable and cause a reduction in the SPMF. However, this picture is not correct for DMO CDM subhalos, and the concept of disruption can conflate three distinct scenarios. Colloquially, a subhalo “disrupting” can mean any of the following:
@@
At first blush, meanings 1 and 2 may seem reasonable or even synonymous. Early CDM simulations and analytic arguments seemed to favor a picture in which subhalos could truly disrupt (e.g. Gonzalez-Casado et al., 1994; Hayashi et al., 2003). These arguments generally rely on comparing the characteristic energy injected into the subhalo by tidal shocks at pericenter to the total binding energy of the subhalo or by considering the properties of an isolated halo profile that was suddenly truncated. Although in detail these arguments break down (van den Bosch et al., 2018), even in modern simulations, individual subhalos can suddenly disappear from subhalo catalogs at high resolutions or can be seen to experience runaway mass loss while they are still resolved by the subhalo finder (e.g., Fig. 16 and Fig. 5 in Mansfield et al. 2023, respectively). But this apparent “disruption” is illusory for most subhalos. While it is true that high-mass subhalos experience so much dynamical friction that they sink to the centers of their host within a few orbits (often remaining self-bound even after the merger finishes: Han et al. 2016; Diemer et al. 2023; Mansfield et al. 2023), lower mass subhalos experience much less dynamical friction (e.g., van den Bosch et al., 2016). Idealized, high-resolution simulations show that low-mass subhalos with the high-internal-density “cuspy” profiles favored by CDM instead stay bound for much longer than the lifetime of the universe (Peñarrubia et al., 2010; van den Bosch & Ogiya, 2018; van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021). This is true even for simulations that include embedded disk potentials (Green et al., 2022). The loss of high-resolution subhalos from simulation catalogs can be explicitly shown to be a combination of subhalo finder failures and numerical non-convergence (Mansfield et al., 2023), while being highly dependent on resolution and subhalo finder.
@@
The near limitless durability of DMO subhalos in $\Lambda$CDM is caused by their density profiles; models that lower the inner density — for example, by modifying the nature of the dark matter particle — can accelerate mass loss rates and even lead to true, resolved unbinding (Nadler et al., 2021; Errani et al., 2023b; Yang et al., 2023; Nadler et al., 2023a; Du et al., 2024). This distinction makes it even more important to be careful about the term “disruption” in CDM contexts, as the term is a correct description of a physical process that is a hallmark of some alternative dark matter models. That said, even for non-$\Lambda$CDM models, it is worth being careful about what is and is not true disruption. The same processes that can numerically destroy DMO $\Lambda$CDM subhalos can accelerate the destruction of subhalos in other frameworks.
@@
The third meaning above is on somewhat better theoretical footing for CDM subhalos: satellite galaxies can always lose enough mass to pass below a given optical observation limit. Studies of galaxy mass loss in hydrodynamic simulations suggest that for a given subhalo, there is a characteristic $M_{\rm vir}/M_{\rm peak}$ scale at which the subhalo’s tidal radius begins to encroach on the satellite’s stars and thus its stellar mass loss rate suddenly accelerates (Smith et al., 2016). But even this meaning has some conceptual and logistical issues. Conceptually, the existence of a characteristic scale in the mass-loss rate does not necessarily imply disruption, just that there is essentially no stellar mass loss soon after infall. Logistically, adopting a star-based approach means that one needs to have a model for the energy/radial distribution of stars in a subhalo, as the end state of a subhalo’s satellite galaxy varies drastically based on these quantities (e.g. Errani et al., 2023a). Additionally, numerical problems associated with the setup of most hydrodynamic simulations mean that they resolve the mass loss of satellite galaxies far more poorly than the disruption of subhalos (Ludlow et al., 2019, 2020, 2021, 2023). This is not an intractable problem, but it does complicate the issue to the point where it is out-of-scope for this work.
@@ 3.2 Interpreting EDEN Results
This distinction holds even when we compare the subhalo peak mass function between DMO and Disk runs (Section 4.2). The peak halo mass $M_{\rm peak}$ is the peak historical value of the halo virial mass $M_{\rm vir}$, which reflects the pre-infall mass of subhalos before experiencing significant tidal stripping from their hosts. In the case of a suppressed SPMF, the increased mass loss rate causes subhalos to approach the numerical-driven limits of the sample more quickly. Given the discussion in Section 3.1, one may question the wisdom in evaluating SPMF at all: after all, at low masses, the SPMF is just the infall mass function minus subhalos that have either numerically disrupted or have been lost by the subhalo finder. Skepticism is warranted, but $M_{\rm peak}$-based measurements serve a legitimate purpose. The fact that subhalos must lose large amounts of dark matter mass prior to losing any stellar mass means that infall- and peak-based mass definitions tend to be better predictors of galaxy properties than the current halo mass (e.g., Reddick et al., 2013). In particular, $M_{\rm peak}$-selected subhalo samples have much higher small-radius number densities around their hosts than $M_{\rm vir}$-selected samples (see review in Section 5.2 of Mansfield et al., 2023), in line with observations.
Indeed, even when analyzed with Symfind, which is capable of tracing $\approx 90\%$ of the subhalos at our minimum mass threshold to the mass scales at which galaxy disruption is likely to begin (Mansfield et al., 2023), the median value of $M_{\rm vir}/M_{\rm peak}$ in our sample is approximately 0.2 in both SymphonyMilkyWay and EDEN. This $M_{\rm vir}/M_{\rm peak}$ threshold is a factor of $2$ to $4$ higher than the masses that hydrodynamic simulations (Smith et al., 2016) and empirical models (Moster et al., 2018; Behroozi et al., 2019; Wang et al., 2024) predict that stellar mass loss should occur. Thus, the vast majority of subhalos considered in this paper would likely still host visible galaxies (if they had one to begin with). This means that the differences in $M_{\rm peak}$ functions discussed below would likely qualitatively extend to stellar mass-selected samples, even if the quantitative translation would require significant modeling work (see Section 5.1 for further discussion).
@@
Figure 5: Subhalo count ratios between EDEN Disk and Symphony DMO simulations for resolved subhalos with $M_{\rm peak}\geqslant 300m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The top panel shows only subhalos within 100 kpc of their hosts, while the bottom panel shows all subhalos within the virial radii of their hosts. Each colored circle is a pair of Symphony/EDEN hosts, with $M_{\ast,\rm Disk}/M_{\rm vir,host}$ values from Symphony. The nine MW-mass hosts we re-simulated with $M_{\ast,\rm Disk}\times 2.5$ are shown with diamonds and lines connecting them with their EDEN fiducial runs. The color bar indicates the stellar half-mass formation scale factor. The Spearman correlation coefficient between $N_{\rm Disk}/N_{\rm DMO}$ and $\log_{10}(M_{\ast,\rm Disk}/M_{\rm vir,host})$ for EDEN fiducial is $r_{S}=-0.60$ for $r<100$ kpc and $r_{S}=-0.74$ for $r<R_{\rm vir,host}$, which are statistically significant negative correlations. We also provide exponential fits to these relations whose best-fit parameters are shown in Table 1. The Spearman $r_{S}$ and exponential fit are carried out on the combined set of EDEN fiducial and EDEN $M_{\ast,\rm Disk}\times 2.5$ simulations. The cyan crosses are the expected number ratio of subhalos in the absence of disk-induced mass loss due to adiabatic contraction (A.C.) from the disk potential. The vertical red (MW, Cautun et al. 2020) and magenta (M31, Watkins et al. 2010; Tamm et al. 2012) bands mark the observational constraints for MW/M31. Orange crosses in the top panel denote the two DMO/Embedded-disk halo pairs from Garrison-Kimmel et al. (2017) whose hydrodynamic counterparts were simulated with FIRE-2 (Wetzel et al., 2023). Disk effects increasingly dominate over adiabatic contraction with increasing $M_{\ast,\rm Disk}/M_{\rm vir,host}$.
@@ 4.1 Disk-to-halo mass ratio as a critical factor for subhalo abundance
At low disk mass, $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$, most EDEN halos demonstrate have $\lesssim 20\%$ fewer subhalos than their disk-less counterparts within 100 kpc and $\lesssim 10\%$ within $R_{\rm vir,host}$. There are a few halos with $N_{\rm Disk}/N_{\rm DMO}>1$. Physically, the addition of the disk can cause the adiabatic contraction of dark matter (Blumenthal et al., 1986; Ryden & Gunn, 1987; Gnedin et al., 2004) and draw more subhalos into the host. Because the disk has two effects, one which increases subhalo abundance and another that decreases it, it is possible that if a low-mass disk’s effect on subhalo mass loss is weak, its host might end up having $N_{\rm Disk}/N_{\rm DMO}>1$.
@@
$M_{\rm DM}(<r_{0})r_{0}=\left[M_{\rm DM}(<r_{0})+M_{\rm Disk}(<r_{f})\right]r_{f}.$ (4)
Here $r_{0}$ is the distance to the MW host center in the DMO run, $M_{\rm DM}(<r_{0})$ is the enclosed dark matter mass within $r_{0}$, which is adiabatically contracted to a smaller sphere with radius $r_{f}$ after adding the disk potential. We assume that the location of the subhalos contract in the same manner as the underlying dark matter density profile of their host, such that subhalos located at $r_{0}$ in SymphonyMilkyWay are expected to be at $r_{f}$. We show the contracted subhalo count ratios within 100 kpc and $R_{\rm vir,host}$ in Fig. 5, which increases with increasing $M_{\ast,\rm Disk}/M_{\rm vir,host}$ as expected. The EDEN hosts having $N_{\rm Disk}/N_{\rm DMO}>1$ within 100 kpc all have low disk-to-halo mass ratios and their subhalo abundance increase are consistent with the predictions of the disk drawing more subhalos inwards due to adiabatic contraction.
At the massive disk mass end, where $M_{\ast,\rm Disk}/M_{\rm vir,host}\gtrsim 0.02$, $N_{\rm Disk}/N_{\rm DMO}\ll 1$ and disk effects dominate over adiabatic contraction. The two EDEN halos with heavy disks that fall within the observational errors of the MW (blue band, Cautun et al. 2020), have $30\sim 40\%$ fewer subhalos within 100 kpc and $20\sim 30\%$ fewer subhalos within $R_{\rm vir,host}$. We also show the subhalo count ratios for the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ and connect them with EDEN fiducial values. The average subhalo abundance suppression in the nine halos with heavier disks are more prominent in both radial ranges (side histograms), all having lower $N_{\rm Disk}/N_{\rm DMO}$ with increasing $M_{\ast,\rm Disk}/M_{\rm vir,host}$. Three of these EDEN $M_{\ast,\rm Disk}\times 2.5$ hosts fall within the M31 disk-to-halo mass ratio range and two of them produces $\gtrsim 50\%$ subhalo abundance suppression within 100 kpc. The one outlier that shows a slight decreased subhalo counts at $r<100$ kpc but an increased counts at $r<R_{\rm vir,host}$ after inserting a $M_{\ast,\rm Disk}\times 2.5$ disk is undergoing a major merger and experiences a boost in outskirt subhalo counts due to its infalling companion.
@@
To summarize our findings, disk effects on subhalo abundance is sensitive to the disk-to-halo mass ratio of the host. The MW and M31 are amongst the most efficient systems in subhalo abundance suppression due to their atypically large $M_{\ast,\rm Disk}/M_{\rm vir,host}$. Disk effect on subhalo abundances is weak for small $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$ hosts. For large $M_{\ast,\rm Disk}/M_{\rm vir,host}\gtrsim 0.02$ hosts, disk effects dominate over adiabatic contraction and become stronger with increasing disk-to-halo mass ratios. We fit the $\log_{10}(M_{\ast,\rm Disk}/M_{\rm vir,host})$-$N_{\rm Disk}/N_{\rm DMO}$ relation jointly for the 45 EDEN fiducial halos and nine EDEN $M_{\ast,\rm Disk}\times 2.5$ halos, using an exponential functional form whose best-fit parameters are provided in Table 1. The Spearman rank correlation coefficients for these joint relations are $r_{S}=-0.60$ within 100 kpc and $r_{S}=-0.74$ within $R_{\rm vir,host}$, both being statistically significant negative correlations between subhalo counts suppression and $M_{\ast,\rm Disk}/M_{\rm vir,host}$.

### v3

@@ 3.1 Background
The language that the field uses to describe mass loss in subhalos paints a picture in which subhalos are completely destroyed in a cataclysmic, discrete event: subhalos are said to “disrupt.” In previous embedded disk simulations (Garrison-Kimmel et al., 2017; Kelley et al., 2019), the abundance suppression in the subhalo peak mass function (SPMF) is used as a diagnostic for ‘disruption’ of subhalos instead of enhanced mass loss. The peak mass of subhalos does not change after infall, so if the subhalos only experience enhanced stripping and could still be tracked by the halo finder, there would be no suppression in the SPMF in the ideal case; whereas ‘disruption’ of subhalos would lead to objects no longer detectable and cause a reduction in the SPMF. However, this picture is not correct for DMO CDM subhalos, and the concept of disruption can conflate three distinct scenarios. Colloquially, a subhalo “disrupting” can mean any of the following:
@@
At first blush, meanings 1 and 2 may seem reasonable or even synonymous. Early CDM simulations and analytic arguments seemed to favor a picture in which subhalos could truly disrupt (e.g. Gonzalez-Casado et al., 1994; Hayashi et al., 2003). These arguments generally rely on comparing the characteristic energy injected into the subhalo by tidal shocks at pericenter to the total binding energy of the subhalo or by considering the properties of an isolated halo profile that was suddenly truncated. Although in detail these arguments break down (van den Bosch et al., 2018), even in modern simulations, individual subhalos can suddenly disappear from subhalo catalogs at high resolutions or can be seen to experience runaway mass loss while they are still resolved by the subhalo finder (e.g., Fig. 16 and Fig. 5 in Mansfield et al. 2024, respectively). But this apparent “disruption” is illusory for most subhalos. While it is true that high-mass subhalos experience so much dynamical friction that they sink to the centers of their host within a few orbits (often remaining self-bound even after the merger finishes: Han et al. 2016; Diemer et al. 2024; Mansfield et al. 2024), lower mass subhalos experience much less dynamical friction (e.g., van den Bosch et al., 2016). Idealized, high-resolution simulations show that low-mass subhalos with the high-internal-density “cuspy” profiles favored by CDM instead stay bound for much longer than the lifetime of the universe (Peñarrubia et al., 2010; van den Bosch & Ogiya, 2018; van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021). This is true even for simulations that include embedded disk potentials (Green et al., 2022). The loss of high-resolution subhalos from simulation catalogs can be explicitly shown to be a combination of subhalo finder failures and numerical non-convergence (Mansfield et al., 2024), while being highly dependent on resolution and subhalo finder.
@@
The near limitless durability of DMO subhalos in $\Lambda$CDM is caused by their density profiles; models that lower the inner density — for example, by modifying the nature of the dark matter particle — can accelerate mass loss rates and even lead to true, resolved unbinding (Nadler et al., 2021; Errani et al., 2023; Yang et al., 2023; Nadler et al., 2023a; Du et al., 2024). This distinction makes it even more important to be careful about the term “disruption” in CDM contexts, as the term is a correct description of a physical process that is a hallmark of some alternative dark matter models. That said, even for non-$\Lambda$CDM models, it is worth being careful about what is and is not true disruption. The same processes that can numerically destroy DMO $\Lambda$CDM subhalos can accelerate the destruction of subhalos in other frameworks.
@@
The third meaning above is on somewhat better theoretical footing for CDM subhalos: satellite galaxies can always lose enough mass to pass below a given optical observation limit. Studies of galaxy mass loss in hydrodynamic simulations suggest that for a given subhalo, there is a characteristic $M_{\rm vir}/M_{\rm peak}$ scale at which the subhalo’s tidal radius begins to encroach on the satellite’s stars and thus its stellar mass loss rate suddenly accelerates (Smith et al., 2016). But even this meaning has some conceptual and logistical issues. Conceptually, the existence of a characteristic scale in the mass-loss rate does not necessarily imply disruption, just that there is essentially no stellar mass loss soon after infall. Logistically, adopting a star-based approach means that one needs to have a model for the energy/radial distribution of stars in a subhalo, as the end state of a subhalo’s satellite galaxy varies drastically based on these quantities (e.g. Errani et al., 2024). Additionally, numerical problems associated with the setup of most hydrodynamic simulations mean that they resolve the mass loss of satellite galaxies far more poorly than the disruption of subhalos (Ludlow et al., 2019b, 2020, 2021, 2023). This is not an intractable problem, but it does complicate the issue to the point where it is out of scope for this work.
@@ 3.2 Interpreting EDEN Results
This distinction holds even when we compare the subhalo peak mass function between DMO and Disk runs (Section 4.2). The peak halo mass $M_{\rm peak}$ is the peak historical value of the halo virial mass $M_{\rm vir}$, which reflects the pre-infall mass of subhalos before experiencing significant tidal stripping from their hosts. In the case of a suppressed SPMF, the increased mass loss rate causes subhalos to approach the numerical-driven limits of the sample more quickly. Given the discussion in Section 3.1, one may question the wisdom in evaluating SPMF at all: after all, at low masses, the SPMF is just the infall mass function minus subhalos that have either numerically disrupted or have been lost by the subhalo finder. Skepticism is warranted, but $M_{\rm peak}$-based measurements serve a legitimate purpose. The fact that subhalos must lose large amounts of dark matter mass prior to losing any stellar mass means that infall- and peak-based mass definitions tend to be better predictors of galaxy properties than the current halo mass (e.g., Reddick et al., 2013). In particular, $M_{\rm peak}$-selected subhalo samples have much higher small-radius number densities around their hosts than $M_{\rm vir}$-selected samples (see review in Section 5.2 of Mansfield et al., 2024), in line with observations.
Indeed, even when analyzed with Symfind, which is capable of tracing $\approx 90\%$ of the subhalos at our minimum mass threshold to the mass scales at which galaxy disruption is likely to begin (Mansfield et al., 2024), the median value of $M_{\rm vir}/M_{\rm peak}$ in our sample is approximately 0.2 in both SymphonyMilkyWay and EDEN. This $M_{\rm vir}/M_{\rm peak}$ threshold is a factor of $2$ to $4$ higher than the masses that hydrodynamic simulations (Smith et al., 2016) and empirical models (Moster et al., 2018; Behroozi et al., 2019; Wang et al., 2024) predict that stellar mass loss should occur. Thus, the vast majority of subhalos considered in this paper would likely still host visible galaxies (if they had one to begin with). This means that the differences in $M_{\rm peak}$ functions discussed below would likely qualitatively extend to stellar mass-selected samples, even if the quantitative translation would require significant modeling work (see Section 5.1 for further discussion).
@@
Figure 5: Subhalo count ratios between EDEN Disk and Symphony DMO simulations for resolved subhalos with $M_{\rm peak}\geqslant 300m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The top panel shows only subhalos within 100 kpc of their hosts, while the bottom panel shows all subhalos within the virial radii of their hosts. Each colored circle is a pair of Symphony/EDEN hosts, with $M_{\ast,\rm Disk}/M_{\rm vir,host}$ values from Symphony. The nine MW-mass hosts we re-simulated with $M_{\ast,\rm Disk}\times 2.5$ are shown with diamonds and lines connecting them with their EDEN fiducial runs. The color bar indicates the stellar half-mass formation scale factor. The Spearman correlation coefficient between $N_{\rm Disk}/N_{\rm DMO}$ and $\log_{10}(M_{\ast,\rm Disk}/M_{\rm vir,host})$ for EDEN fiducial is $r_{S}=-0.60$ for $r<100$ kpc and $r_{S}=-0.74$ for $r<R_{\rm vir,host}$, which are statistically significant negative correlations. We also provide exponential fits to these relations whose best-fit parameters are shown in Table 1. The Spearman $r_{S}$ and exponential fit are carried out on the combined set of EDEN fiducial and EDEN $M_{\ast,\rm Disk}\times 2.5$ simulations. The cyan crosses are the expected number ratio of subhalos in the absence of disk-induced mass loss due to adiabatic contraction (A.C.) from the disk potential. The vertical red (MW, Cautun et al. 2020) and magenta (M31, Watkins et al. 2010; Tamm et al. 2012) bands mark the observational constraints for MW/M31. Orange crosses in the top panel denote the two DMO/Embedded-disk halo pairs from Garrison-Kimmel et al. (2017) whose hydrodynamic counterparts were simulated with FIRE-2 (Wetzel et al., 2023). Disk stripping dominates over adiabatic contraction across the range of $M_{\ast,\rm Disk}/M_{\rm vir,host}$ covered by EDEN.
@@ 4.1 Disk-to-halo mass ratio as a critical factor for subhalo abundance
At low disk mass, $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$, most EDEN halos demonstrate $\lesssim 20\%$ fewer subhalos than their disk-less counterparts within 100 kpc and $\lesssim 10\%$ within $R_{\rm vir,host}$. There are a few halos with $N_{\rm Disk}/N_{\rm DMO}>1$. Physically, the addition of the disk can cause the adiabatic contraction of dark matter (Blumenthal et al., 1986; Ryden & Gunn, 1987; Gnedin et al., 2004) and draw more subhalos into the host. Because the disk has two effects, one which increases subhalo abundance and another that decreases it, it is possible that if a low-mass disk’s effect on subhalo mass loss is weak, its host might end up having $N_{\rm Disk}/N_{\rm DMO}>1$.
@@
$\left(1+\frac{M_{\ast,\rm Disk}}{M_{\rm vir,host}}\right)M_{\rm DM}(<r_{0})r_{0}=\left[M_{\rm DM}(<r_{0})+M_{\rm Disk}(<r_{f})\right]r_{f}.$ (4)
Here $r_{0}$ is the distance to the MW host center in the DMO run, $M_{\rm DM}(<r_{0})$ is the enclosed dark matter mass within $r_{0}$, which is adiabatically contracted to a smaller sphere with radius $r_{f}$ after adding the disk potential. The mass component of the disk potential before contraction is assumed to follow that of the dark matter in the DMO run. We assume that the locations of the subhalos contract in the same manner as the underlying dark matter density profile of their host, such that subhalos located at $r_{0}$ in SymphonyMilkyWay are expected to be at $r_{f}$. We show the contracted subhalo count ratios within 100 kpc and $R_{\rm vir,host}$ in Fig. 5, which creates a $\sim 5\%$ subhalo count boost within 100 kpc and no effect out to $R_{\rm vir,host}$. The EDEN hosts having $N_{\rm Disk}/N_{\rm DMO}>1$ within 100 kpc all have low disk-to-halo mass ratios and their subhalo abundance increase is consistent with the predictions of adiabatic contraction.
At the massive disk mass end, where $M_{\ast,\rm Disk}/M_{\rm vir,host}\gtrsim 0.02$, $N_{\rm Disk}/N_{\rm DMO}\ll 1$ and disk stripping effects dominate over adiabatic contraction. The two EDEN halos with heavy disks that fall within the observational errors of the MW (blue band, Cautun et al. 2020), have $30\sim 40\%$ fewer subhalos within 100 kpc and $20\sim 30\%$ fewer subhalos within $R_{\rm vir,host}$. We also show the subhalo count ratios for the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ and connect them with EDEN fiducial values. Three of the $M_{\ast,\rm Disk}\times 2.5$ hosts have $M_{\ast,\rm Disk}/M_{\rm vir,host}$ within the observational uncertainties of M31 (Watkins et al., 2010; Tamm et al., 2012) and show even larger suppression on average. The average subhalo abundance suppression in the nine halos with heavier disks is more prominent in both radial ranges (side histograms), all having lower $N_{\rm Disk}/N_{\rm DMO}$ with increasing $M_{\ast,\rm Disk}/M_{\rm vir,host}$.
Three of these EDEN $M_{\ast,\rm Disk}\times 2.5$ hosts fall within the M31 disk-to-halo mass ratio range and two of them produce $\gtrsim 50\%$ subhalo abundance suppression within 100 kpc. The one outlier (Halo 675) that shows a slight decrease in subhalo counts at $r<100$ kpc but an increase in count at $r<R_{\rm vir,host}$ after inserting a $M_{\ast,\rm Disk}\times 2.5$ disk is undergoing a major merger and experiences a boost in outskirt subhalo counts due to its infalling companion.
@@
To summarize our findings, disk effects on subhalo abundance are sensitive to the disk-to-halo mass ratio of the host. The MW and M31 are amongst the most efficient systems in subhalo abundance suppression due to their atypically large $M_{\ast,\rm Disk}/M_{\rm vir,host}$. Disk effects on subhalo abundances are weak for small $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$ hosts. For large $M_{\ast,\rm Disk}/M_{\rm vir,host}\gtrsim 0.02$ hosts, disk effects dominate over adiabatic contraction and become stronger with increasing disk-to-halo mass ratios. We fit the $\log_{10}(M_{\ast,\rm Disk}/M_{\rm vir,host})$-$N_{\rm Disk}/N_{\rm DMO}$ relation jointly for the 45 EDEN fiducial halos and nine EDEN $M_{\ast,\rm Disk}\times 2.5$ halos, using an exponential functional form whose best-fit parameters are provided in Table 1. The Spearman rank correlation coefficients for these joint relations are $r_{S}=-0.60$ within 100 kpc and $r_{S}=-0.74$ within $R_{\rm vir,host}$, both being statistically significant negative correlations between subhalo count suppression and $M_{\ast,\rm Disk}/M_{\rm vir,host}$.